In this lab exercise a solution containing K+ is passed down an ion exchange column containing a cation exchange resin in the acid form. The K+ ions in the solution replace the resin bound H3O+ ions which elute from the column and then are titrated. This means that the column now holds the K+ ions. To use the column again we must "recharge" the column to remove K+ ions. In your Lab Notebook, describe the exact procedure you will follow to "recharge" the exchange column for the next use.

A 0.171 gram solution of KxFe(C2O4)yzH2O is dissolved in water and the solution is passed down a column containing a cation exchange resin in the acid form. The eluted solution is then titrated with a standard NaOH solution to the first equivalence point. Experimentally it required 10.65 mL of a 0.132 M NaOH solution to reach the first equivalence point.

Calculate the number of moles of K+ in the weighted sample.

Calculate the mass (grams) of K+ in the weighted sample.

Calculate percent of potassium in the weighted sample.
Answer #1

(a) Given : concentration of NaOH = 0.132 M

volume of NaOH solution used = 10.65 mL = 0.01065 L

moles of NaOH used = (concentration of NaOH) * (volume of NaOH solution used)

moles of NaOH used = (0.132 M) * (0.01065 L)

moles of NaOH used = 1.406 x 10-3 mol

moles of H3O+ in eluted solution = moles of NaOH used

moles of H3O+ in eluted solution = 1.406 x 10-3 mol

moles K+ in sample = moles of H3O+ in eluted solution

moles K+ in sample = 1.406 x 10-3 mol

(b) mass K+ in sample = (moles K+ in sample) * (molar mass K)

mass K+ in sample = (1.406 x 10-3 mol) * (39.1 g/mol)

mass K+ in sample = 0.0550 g

(c) percent of potassium in the weighted sample = (mass K+ in sample / mass of sample) * 100

percent of potassium in the weighted sample = (0.0550 g / 0.171 g) * 100

percent of potassium in the weighted sample = (0.321) * 100

percent of potassium in the weighted sample = 32.1 %